# The Popularity Hypothesis in Software Security: A Large-Scale Replication with PHP Packages

### v1

@@ 4.1 Descriptive Statistics
Regarding multicollinearity, Fig. 1 displays (Pearson’s) correlation coefficients across all metrics in the dataset by using the logarithm transformation. As can be seen, all coefficients have positive signs and some of these are relatively large in their magnitudes. These correlations are nothing surprising as such. Particularly when operating with large behavioral datasets—a domain to which the dataset examined can also be seen to belong, everything tends to be correlated with everything [52]. In any case, stars, watchers, and installs are moderately or even strongly correlated with each other on one hand and installs and releases on the other, to use the thresholds and adjectives for these from the replicated study [1]. Of these correlations, particularly the latter two seem sensible in a sense that more releases may cause more installs. These correlations may affect the classifications because particularly the naïve Bayes and logistic regression rely on an independence assumption. Though, even naïve Bayes seems to still perform well under multicollinearity [53]. Therefore, it is more relevant to continue by pointing out the only modest magnitudes between security and rest of the metrics. This observation might be taken as a prior expectation that the hypothesis may not hold.
@@
However, the descriptive results in Fig. 2 tell a different story. These are based on conventional $t$-test estimates using a correction for unequal variances, which are evident also according to the Leneve’s classical test [54]. As can be seen, all means are different in the two groups separated by the truncated, binary-valued security metric. In other words, both popular packages and packages with long release histories seem to have been affected by vulnerabilities over the years. This observation provides a good motivation to briefly also report regression analysis results before continuing to the classification results.
4.2 Regression Analysis
@@
Regression results can shed a little more light on the earlier correlation results in Fig. 1. Thus, coefficients from three OLS regressions for the three models noted in Subsection 3.2.3 are shown in Fig. 3. The logarithm transformation is again used for all metrics. All coefficients are statistically significant at a 99% confidence level, which is hardly surprising due to the sample size [55]. As could be furthermore expected, the performance increases the more there are metrics. The full model yields an adjusted $R^{2}=0.045$, meaning that roughly about five percent of the total variance is explained by the seven metrics. When keeping in mind that there are only seven hundred seventy seven observations for which security attains a value larger than zero, the performance is not necessarily that bad. Of the coefficients for the individual metrics, the one for suggesters stands out. Interestingly, the coefficient for installs has a negative sign, which seems to contradict the earlier results in Fig. 2. All in all, nevertheless, the evidence is again on the side of supporting the hypothesis. The magnitudes and signs of the coefficients are presumably linked to the correlations between the explanatory metrics.
4.3 Classification
@@
The classification results are summarized in Tables 2, 3, and 4 for the three classifiers. In each table, the second, third, and fourth columns denote the three models estimated. To recall: the first model is the barebone one with just the releases metric, the second model adds the sum variable of the popularity metrics, and the third model uses all seven metrics individually. Then, the rows are arranged according to the three performance metrics described in Subsection 3.2.2. For each metric, the results are reported according to the four balancing solutions.
Table 2: Naïve Bayes Classification Results1
@@
  • 1.
    1 The best values are colored in each of the three rowwise panels.
Table 3: Boosted Logistic Regression Classification Results1
@@
[…]
  • 1.
    1 The best values are colored in each of the three rowwise panels.
Table 4: Random Forest Classification Results1
@@
  • 1.
    1 The best values are colored in each of the three rowwise panels.
With these notes, the dissemination of the classification results can be started by noting that the balancing solutions improve the performance significantly. Although not reported, plain unbalanced classifications yield much worse performance, regardless of the three classifiers and the three models. Regarding the balancing solutions, it seems that the ROSE algorithm is better than the rest, although SMOTE does well with the Naïve Bayes classifier, and plain oversampling yields the best performance in one outlying case. As for the classifiers, the boosted logistic regression outperforms the Naïve Bayes and random forest classifiers according to the BA and G-mean metrics. According to these two performance metrics, the highest value reached is $0.840$, which can be interpreted as a decent value in the extreme imbalance context operated.
@@
5 Conclusion
@@
The paper revisited a popularity hypothesis in empirical software security research; an assumption that a software’s popularity can explain to some degree the software’s insecurity. According to the results presented, the hypothesis can be concluded to hold: popular PHP packages have been affected by vulnerabilities throughout their release histories more often than unpopular PHP packages. By implication, the paper also fails to replicate a previous study’s conclusion, although it should be noted that the study strictly speaking only concluded that popularity is not a reliable indicator of software security [1]. Although a decent classification performance was obtained even with the limited amount of explanatory information used, the results reported should not be taken to mean that popularity alone could explain the security or insecurity of packages written in the PHP programming language.
@@
The confirmation of the hypothesis has implications also for empirical research more broadly. Recently, it has been argued that probabilistic random sampling should be preferred in empirical software engineering to fix a real or perceived generalizability crisis [5]. The results presented cast a doubt upon the argument. The unconditional probability of picking a non-vulnerable PHP package from the almost entire Packagist population observed is $0.998$. In other words, it would be very likely that only non-vulnerable packages would end up in a sample picked randomly. By implication, it is necessary to return to the concept of reported vulnerabilities discussed in the introduction. Thus, should one consequently believe that the $381,216$ packages in the sample without reported vulnerabilities are free of security issues and generally of high quality? Although no definite answers can be given, many would likely prefer a negative answer to the question. If this answer is accepted, the why-question subsequently emerges. As was noted, the famous Linus law probably has something to do also with reporting of vulnerabilities in large software ecosystems. If a package is particularly popular, it may be more likely that some developers actually also review and even audit its code base.
@@
The confirmation of the popularity hypothesis has also other implications. As was noted, popularity metrics have also been used to grant cyber security funding for open source software projects. As the funding grants have involved also testing and security audits [8], it may be that the evidence for the popularity hypothesis also strengthens in the future in a sense that even more vulnerabilities are reported for popular open source software projects. The same point applies with respect to automated tools. For instance, also large-scale security scanning efforts of open source software projects have used popularity and related metrics to select preferable samples for scanning [56]. Audits and scanning are also important topics because they are arguably closer to “real” software security than what is available through observing reported vulnerabilities. To this end, it can be argued that also construct validity should be revisited, discussed, and debated in empirical software security research. In the meanwhile, it suffices to reiterate an argument that reported vulnerability counts should be approached with care when making practical choices about open source software packages [19]. This argument reinforces the earlier remark about long-standing philosophical and theoretical problems in applying principles of empirical sciences to cyber security research.
@@
Finally, both the replicated study and this replication suffer from potential generalizability issues. In other words, it remains unclear whether the hypothesis would, or would not, hold with a further different dataset. As said, however, it is also generally unclear how the generalizability problem should be addressed—and whether it is even possible to address it in empirical software security research and empirical software engineering in general. Regardless, as it stands, there is now conflicting evidence about the popularity hypothesis. It is up to a reader to make the final verdict about whether software’s popularity affects its insecurity in light of the evidence put forward.

### v2

@@ IV-A Descriptive Statistics
Regarding multicollinearity, Fig. 2 displays Pearson’s correlation coefficients across all metrics in the dataset by using the logarithm transformation. As can be seen, all coefficients have positive signs and some of these are relatively large in their magnitudes. These correlations are nothing surprising as such. Particularly when operating with large behavioral datasets—a domain to which the dataset examined can also be seen to belong, everything tends to be correlated with everything [61]. In any case, stars, watchers, and installs are moderately or even strongly correlated with each other on one hand and installs and releases on the other, to use the thresholds and adjectives for these from one of the replicated studies [1]. Of these correlations, particularly the latter two seem sensible in a sense that more releases may cause more installs. Another point worth making is that the observations are in contrast with the other replicated study, which concluded that reported vulnerability counts “show almost no relationship with metrics like stars or forks” [2, p. 4]. Furthermore, these correlations may affect the classifications because particularly the naïve Bayes and logistic regression rely on an independence assumption. Though, even naïve Bayes seems to still perform well under multicollinearity [62]. Therefore, it is more relevant to continue by pointing out the only modest magnitudes between security and rest of the metrics. This observation might be taken as a prior expectation that the hypothesis may not hold.
@@
However, the descriptive results in Fig. 3 tell a different story. These are based on conventional $t$-test estimates using a correction for unequal variances, which are evident also according to the Leneve’s classical test [63]. As can be seen, all means are different in the two groups separated by the truncated, binary-valued security metric. In other words, both popular packages and packages with long release histories seem to have witnessed reported vulnerabilities over the years. This observation provides a good motivation to briefly also report regression analysis results before continuing to the classification results.
IV-B Regression Analysis
@@
Regression results can shed a little more light on the earlier correlation results in Fig. 2. Thus, coefficients from three OLS regressions for the three models noted in Subsection III-B3 are shown in Fig. 4. The logarithm transformation is again used for all metrics. All coefficients are statistically significant at a 99% confidence level, which is hardly surprising due to the sample size [64]. As could be furthermore expected, the performance increases the more there are metrics. The full model yields an adjusted $R^{2}=0.045$, meaning that roughly about five percent of the total variance is explained by the seven metrics. When keeping in mind that there are only seven hundred seventy seven observations for which security attains a value larger than zero, the performance is not necessarily that bad. Of the coefficients for the individual metrics, the one for suggesters stands out. Interestingly, the coefficient for installs has a negative sign, which seems to contradict the earlier results in Fig. 3. All in all, nevertheless, the evidence is again on the side of supporting the hypothesis. The magnitudes and signs of the coefficients are presumably linked to the correlations between the explanatory metrics.
IV-C Classification
@@
The classification results are summarized in Tables II, III, and IV for the three classifiers. In each table, the second, third, and fourth columns denote the three models estimated. To recall: the first model is the barebone one with just the releases metric, the second model adds the sum variable of the popularity metrics, and the third model uses all seven metrics individually. The rows are arranged according to the three performance metrics described earlier. For each metric, the results are reported according to the balancing solutions.
TABLE II: Naïve Bayes Classification Results1
@@
  • 1 The best values are colored in each of the three rowwise panels.
TABLE III: Boosted Logistic Regression Classification Results1
@@
ROSE 0.455 0.425 0.118
  • 1 The best values are colored in each of the three rowwise panels.
TABLE IV: Random Forest Classification Results1
@@
  • 1 The best values are colored in each of the three rowwise panels.
With these notes, the dissemination of the classification results can be started by noting that the balancing solutions improve the performance significantly. Although not reported, plain unbalanced classifications yield much worse performance, regardless of the three classifiers and the three models. Regarding the balancing solutions, it seems that the ROSE algorithm is better than the rest, although SMOTE does well with the Naïve Bayes classifier, and plain oversampling yields the best performance in one outlying case. As for the classifiers, the boosted logistic regression outperforms the Naïve Bayes and random forest classifiers according to the BA and G-mean metrics. According to these two performance metrics, the highest value reached is $0.840$, which can be interpreted as a decent value in the extreme imbalance context.
@@
V Discussion
In what follows, the conclusion is first summarized. A few points about implications are subsequently discussed.
V-A Conclusion
@@
The paper revisited a popularity hypothesis in empirical software security research; an assumption that a software’s popularity can explain to some degree the software’s insecurity. According to the results presented, the hypothesis can be concluded to hold: popular PHP packages have seen more reported vulnerabilities throughout their release histories than unpopular PHP packages. By implication, the paper also fails to replicate a previous study’s conclusion, although it should be noted that the study strictly speaking only concluded that popularity is not a reliable indicator of software security [1]. Although a decent classification performance was obtained even with the limited amount of explanatory information used, the results reported should not be taken to mean that popularity alone could explain the security or insecurity of packages written in the PHP programming language.
V-B Research Implications
@@
The confirmation of the hypothesis has implications also for empirical research more broadly. Recently, it has been argued that probabilistic random sampling should be preferred in empirical software engineering to fix a real or perceived generalizability crisis [9]. Analogously to previously solicited expert opinions [65], the results presented cast a doubt upon the argument. The unconditional probability of picking a PHP package without reported vulnerabilities from the almost entire Packagist population observed is $0.998$. In other words, it would be very likely that only packages without reported vulnerabilities would end up in a sample picked randomly. By implication, it is necessary to return to the concept of reported vulnerabilities discussed in the introduction. Thus, should one consequently believe that the $381,216$ packages in the sample without reported vulnerabilities are free of security issues and generally of high quality? Although no definite answers can be given, many would likely prefer a negative answer to the question. If this answer is accepted, the why-question subsequently emerges. As was noted, the famous Linus law probably has something to do also with reporting of vulnerabilities in large software ecosystems. If a package is especially popular, it may be more likely that some developers actually also review and even audit its code base.
V-C Practical Implications
@@
The confirmation of the popularity hypothesis has also other implications. As was noted, popularity metrics have also been used to grant cyber security funding for open source software projects. As the funding grants have involved also testing and security audits [12], it may be that the evidence for the popularity hypothesis also strengthens in the future in a sense that even more vulnerabilities are reported for popular open source software projects. The same point applies with respect to automated tools. For instance, also large-scale security scanning efforts of open source software projects have used popularity and related metrics to select preferable samples for scanning [66]. Audits and scanning are also important topics because they are arguably closer to “real” software security than what is available through observing reported vulnerabilities. To this end, it can be argued that also construct validity should be revisited, discussed, and debated in empirical software security research. In the meanwhile, it suffices to reiterate an argument that reported vulnerability counts should be approached with care when making practical choices about open source software packages [23]. A related point is that software popularity metrics—like all metrics—can be manipulated. Also this point has security consequences because it has been observed that manipulated popularity metrics have been used on GitHub to push malware repositories [67]. These arguments reinforce the earlier remark about long-standing philosophical and theoretical problems in applying principles of empirical sciences to cyber security research.
V-D Concluding Remarks
Finally, it should be emphasized that the results suggest that popular packages tend to have more reported vulnerabilities but this observation does not mean that popularity causes insecurity. Among other things, it is possible that popular packages are simply more exposed and get more attention from open source software communities and security researchers, which presumably increases the chances of finding and reporting vulnerabilities. Thus, the empirical patterns observed should not be interpreted as an evidence of a causal relationship.
@@ V-D Concluding Remarks
Furthermore, both the replicated studies and this replication suffer from a generalizability problem. That is, it remains unclear whether the hypothesis would, or would not, hold with a further different dataset. As said, however, it is also generally unclear how the problem should be addressed—and whether it is even possible to address it in empirical software security research and empirical software engineering in general. Regardless, as it stands, there is now conflicting evidence about the popularity hypothesis. It is up to a reader (cf. [68, p. 76]) to make the final verdict about whether software’s popularity affects its insecurity in light of the evidence put forward.